eviant:

PLS HELP!!! Find the distance between R(–1, 1) and S(–4, –5) to the nearest tenth.

eviant:

@Vocaloid

Shadow:

\[D = \sqrt (x _{2-} x _{1})^2 + (y _{2} - y _{1})^2\]

eviant:

uh?

eviant:

@bm717 what was that?

Shadow:

Ignore her. Do you know how to use the formula?

eviant:

the distance formula?

Shadow:

Yes

bm717:

It was a hand spasm I had this problem first ;-;

eviant:

I tried doing the distance formula, and I came up with the wrong answer

Shadow:

Can you show your setup?

eviant:

yes, I did (-1-4) to the second power which would be -25

Shadow:

After you take -5 to the second power, the negative sign goes away \[-5^2 = -5 \times -5 = 5 \times 5 = 25\]

eviant:

then I did (1-(-5))to the second power, which would be 36

eviant:

then I added 25+36=61

eviant:

then I did the square root of 61, which is 7.8

Shadow:

I think I see your problem. You need to do x1 - x2 x1 is -1 x2 is -4 -1 - (-4) = -1 + 4 = 3

eviant:

so, it's 3 +36

Shadow:

Yes

Shadow:

Try that now

eviant:

so, square root of 39 is 6.2

eviant:

@Shadow the choices are: 6.7 9.0 15.8 45.0

Shadow:

You forgot to square 3

bm717:

I gave shadow the answers, please medal me k thanks

eviant:

@shadow what?

Shadow:

You are supposed to do 3^2 + 6^2 You did 3 + 6^2

eviant:

so, I should add the square roots of 3 and 6

Shadow:

Do 3^2

eviant:

so, 6.7

Shadow:

Correct